.A punter can kick a football with an initial velocity of 48 feet per second. How many seconds will it take for the ball to return to the ground? (Hint: use the formula h= vt – 16t2.)
Answer by stanbon(75887)

A punter can kick a football with an initial velocity of 48 feet per second. How many seconds will it take for the ball to return to the ground? (Hint: use the formula h= vt – 16t2.)
------------
Height(t) = -16t^2+48t
----
When the ball is on the ground, height is zero.
---------------------
Solve -16t^2+48t = 0
-16t(t-3) = 0
Positive solution:
t = 3 seconds
======================
Cheers,
Stan H.